OpenStudy (anonymous):

Two triangles can be formed with the given information. Use the Law of Sines to solve the triangles. (5 points) A = 56°, a = 12, b = 14 B = 75.3°, C = 48.7°, c = 10.9; B = 104.7°, C = 19.3°, c = 4.8 B = 14.7°, C = 109.3°, c = 13.7; B = 165.3°, C = 70.7°, c = 13.7 B = 75.3°, C = 48.7°, c = 13.2; B = 104.7°, C = 19.3°, c = 13.2 B = 14.7°, C = 109.3°, c = 10.5; B = 165.3°, C = 70.7°, c = 10.5

OpenStudy (anonymous):

@SolomonZelman @ganeshie8 @AlexandervonHumboldt2

OpenStudy (anonymous):

@mathstudent55

OpenStudy (mathstudent55):

Do you know the law of sines?

OpenStudy (anonymous):

I have heard of it only

OpenStudy (anonymous):

it is a/sinA=b/sinB=c/sinC, right?

OpenStudy (anonymous):

Is that right @mathstudent55

OpenStudy (mathstudent55):

Yes. That is it.

OpenStudy (mathstudent55):

The law if sines states that the ratio of the length of a side of a triangle to the sine of the opposite angle is equal for all sides and opposite angles.

OpenStudy (anonymous):

so to find the angle B I need to solve sinB=b*sinA/a

OpenStudy (mathstudent55):

In order to use the law of sines, you need to find that ratio. In order to do that, you must know the length of one side and the measure of the opposite angle to that side.

OpenStudy (anonymous):

sinB=14*sin56^deg/12 it is about 75.3 degrees.

OpenStudy (mathstudent55):

Look at the given info. You are told A = 56 deg, and a = 12 That is what you need to find the ratio. Then since you are given one other side length, you can find the opposite angle.

OpenStudy (anonymous):

c=12*sin48.7deg/sin56deg

OpenStudy (anonymous):

ok, is the correct answer A?

OpenStudy (mathstudent55):

\(\dfrac{a}{\sin A} = \dfrac{b}{\sin B} = \dfrac{c}{\sin C} \) \(\dfrac{a}{\sin A} = \dfrac{b}{\sin B} \) \(\dfrac{12}{\sin 56^o} = \dfrac{14}{\sin B} \) \(\sin B = \dfrac{14\sin 56^o}{12} \) \(B = \sin^{-1} \dfrac{14\sin 56^o}{12} \) \(B \approx 75.3^o\) You are correct.

OpenStudy (mathstudent55):

Now you find C: C = 180 - 75.3 - 56 = 48.7 deg and you do the law of sines for c and C

OpenStudy (mathstudent55):

\(\dfrac{a}{\sin A} = \dfrac{c}{\sin C}\) \(\dfrac{12}{\sin 56^o} = \dfrac{c}{\sin 48.7^o}\) \(c = \dfrac{12\sin 48.7^o}{\sin 56^o} \) \(c = 10.9\)

OpenStudy (mathstudent55):

Yes, the answer is A.

OpenStudy (mathstudent55):



OpenStudy (mathstudent55):